# Confidence-based Estimators for Predictive Performance in Model Monitoring

### v1

@@ 4 Average Confidence
In this section, we present some theoretical results for the AC method, which was originally proposed as a baseline method for detecting misclassified and out-of-distribution instances in neural networks (Hendrycks and Gimpel, 2017). When used in unsupervised accuracy estimation, the predictive accuracy is estimated to be the average of confidence scores, hence the name.
@@
Since $|F_{k}|=\binom{n}{k}$, calculating the required probabilities using the definition directly is infeasible even for relatively small values of $n$. Fortunately, there are several ways to avoid this combinatorial explosion. In this work, we leverage the algorithm presented in Hong (2013), which is based on the Fast Fourier Transform and implemented in the poibin [^1] https://github.com/tsakim/poibin python library. It allows us to derive the Poisson binomial distributions exactly and almost instantly.
@@ 4.3 Estimating the confusion matrix for failure prediction
Sometimes confidence scores are used as a tool for failure prediction (Hendrycks and Gimpel, 2017; Kivimäki et al., 2023). A failure prediction mechanism based on confidence scores is a mapping $h:S\rightarrow\{0,1\}$. It sets a confidence threshold $t\in[0,1]$ and treats all predictions with a confidence score equal to or above the threshold as correct and below the threshold as incorrect. Predictions with sub-threshold confidence can be altogether discarded, or sent to further inspection by some downstream system or human-in-the-loop. This might be useful, especially in situations where different types of errors induce different costs. It is worthwhile to stress that this treatment creates second-order predictions (predictions about predictions) and is not to be confused with the true underlying discriminator $g$.
@@ 5 Experiments
We conduct two experiments with synthetic data. In the first experiment, we examine the quality of point predictions and confidence intervals of the estimates derived using the AC method when the confidence scores are known to be calibrated. In the second experiment, we compare AC, DoC-Feat, and ATC under covariate shift when used to estimate predictive accuracy in a monitoring setting.
@@ 5.1.1 Data description
We chose to create our simulated set of confidence scores by drawing samples from a mixture of three Beta distributions. The first component was biased to generate high-confidence predictions, the second to generate average ones, and the third to create low-confidence predictions. By adjusting the mixture weights, we could effectively emulate distributions of confidence scores outputted by models operating on real-life data. Then, by changing the mixture weights, we could easily simulate shifts in model performance to see how the estimator under scrutiny managed to keep track of them. The original dataset was trying to emulate a typical distribution of confidence scores where most predictions are made with high confidence (Guo et al., 2017). For the shifted distribution, the fraction of high-confidence predictions was slightly decreased. The exact parameters and weights of the mixture components used in the experiment are given in Table 2.
Table 2: Parameters and weights of the Beta mixture components
@@ 5.1.2 Quality of point estimates under shift
We tested the quality of point estimates for predictive accuracy derived using AC by simulating gradual data shifts. We used two different monitoring window sizes, namely 100 and 500 predictions. We gradually increased the fraction of instances sampled from the shifted distributions in increments of $5\%$, starting from $0\%$ and ending in $100\%$. For each degree of shift, we conducted 1,000 trials, where we measured the error of the estimate in each of the trials and plotted the mean of the errors in Fig. 2, along with one standard deviation (shaded area). We contrasted the results of AC by also similarly plotting estimates derived by using either ATC (Garg et al., 2022) or a binomial assumption (using the overall accuracy of the original distribution as the parameter). The binomial assumption comparison was included to give a sense of how fast the estimation error would increase if the predictive performance was assumed not to change.
@@
Figure 2: Left: Distribution of confidence scores in the simulated data. Right: The quality of point estimates under gradual data drift.
@@ 5.1.3 Quality of confidence intervals
where $\sigma$ is an estimation of the population variance as the variance of the true labels in the reference data set (where GT is known) and $n$ is the monitoring window size (NannyML uses the term "chunk size"). We replaced the ±$\pm 3$ standard errors with ±$\pm 1.96$ standard errors in our tests to get the lower and upper $95\%$ confidence limits for a fair comparison against the CIs derived using the Poisson binomial approach. The results of this experiment are presented in Fig. 3, clearly showing the superiority of the Poisson binomial approach, which produces (roughly) valid CIs. In contrast, the confidence bands produced by the SEM approach are extremely conservative.
@@
We argue that it is conceptually clearer to account for the variance due to sampling effects with control limits and form the confidence bands for the estimated accuracy solely on the basis of uncertainty in the estimates themselves using the Poisson binomial approach. In this way, all available information from each batch gets utilized and the confidence bands are allowed to vary from batch to batch. We are currently working to implement this improved approach to forming confidence bands into the NannyML library.
@@
5.2 Experimenting under covariate shift
Next, we compared the confidence-based accuracy estimators AC, DoC-feat, and ACT with both uncalibrated and calibrated confidence scores. Earlier studies (Guillory et al., 2021; Garg et al., 2022) have had conflicting results on whether calibration has a beneficial effect in estimating predictive accuracy or not. In these studies, accuracy estimation has been performed on the dataset level (with sample size in the order of thousands or tens of thousands), whereas we are interested in estimates derived for sample sizes generally used in model monitoring (in the order of hundreds). Furthermore, they have emphasized neural networks used in image classification, with the nature of shifts in the shifted datasets not fully controlled. We expanded this scope by comparing the estimation methods using seven non-neural classifier models, namely Logistic Regression (LR), (Gaussian) Naïve Bayes (NB), K-Nearest Neighbors (KNN), Support Vector Machine (SVM) with RBF kernel, and Random Forest (RF), using implementations provided by the scikit-learn library  [^3] https://scikit-learn.org/stable/, as well as XGBoost (XGB)  [^4] https://xgboost.readthedocs.io/en/stable/, and Light-GBM (LGBM)  [^5] https://lightgbm.readthedocs.io/en/latest/index.html to see whether the earlier results would generalize to these models as well. All models were trained with the default parameter settings of their respective implementations.
@@ 5.2.1 Data and model training descriptions
We used simulated data in order to have full control over the nature of the shift. Since all confidence-based accuracy estimators will eventually fail under concept shift (Garg et al., 2022), we restricted the shift as covariate shift. We experimented with two scenarios, where in both scenarios our data consisted of two features and a binary label (1 or 0).
@@
In both scenarios, we trained the models with 100,000 samples, of which $80\%$ were easy and $20\%$ were hard. Using the same ratio, we drew 25,000 additional samples to train the calibration mappings for each model and 25,000 additional samples for the setup required by the DOC-Feat and ATC methods. Since there was no shortage of data, we used the non-parametric Isotonic Regression (Zadrozny and Elkan, 2002) to derive the calibration mappings. It is generally considered to produce good quality calibration mappings but is also known to sometimes overfit with smaller datasets (Kull et al., 2017). In each trial, we created a test dataset of 25,000 samples, with an increasing portion of hard-to-predict samples, as described in Table 3.
Table 3: The number of easy and hard samples for each degree of shift
@@ 5.2.2 Results with linear decision boundary
Table 5: Mean absolute error of estimated accuracy using data with linear decision boundary. Underscipts $u$ and $c$ refer to using uncalibrated or calibrated confidence scores respectively. Calibration errors are calculated on the dataset level.
@@ 5.2.3 Results with non-linear decision boundary
Table 7: Mean absolute error of estimated accuracy using data with non-linear decision boundary. Underscipts $u$ and $c$ refer to using uncalibrated or calibrated confidence scores respectively. Calibration errors are calculated on the dataset level.
@@ 6.1 Novel contributions
Earlier studies (e.g. (Guillory et al., 2021; Garg et al., 2022; Deng et al., 2023; Baek et al., 2022; Chen et al., 2021a; Lu et al., 2023)) have explored the problem of unsupervised accuracy estimation only on the dataset level, that is, whether the suggested methods would be able to correctly estimate the accuracy of an ML model on some given shifted dataset, in the absence of GT labels. Furthermore, theoretical guarantees for the proposed methods are provided only by Chen et al. (2021a) and Garg et al. (2022). In real-life model monitoring, instead of complete, shifted datasets, users typically have access only to relatively small samples of the potentially shifted data and are mostly interested in detecting relevant changes in the underlying data distribution on the fly to trigger appropriate interventions. In this study, we have tried to bridge this gap by exploring different confidence-based performance estimators in a monitoring setting.
@@ 6.2 Remarks on AC and other confidence-based estimators
Ever since AC was introduced as a method for OoD-detection (Hendrycks and Gimpel, 2017), it has become a de facto baseline for confidence-based estimators. However, up until now, there has never been a proper justification for its use as such or a thorough analysis of its properties. As confidence-based estimators (such as CBPE in NannyML) are already being implemented in monitoring libraries intended for industrial use, such theoretical analysis is long overdue. In this paper, we have provided such a justification by showing that under the calibration assumption, AC is an unbiased and consistent estimator of model accuracy. Another desirable property of AC is that one can leverage the properties of the Poisson binomial distribution to derive CIs for the estimates, which are shown to be valid under the same assumption. None of the other methods have been reported to have this property.
Although using control limits and CIs to account for sampling effects and uncertainty in the estimates respectively, two other sources of uncertainty remain. First, in this work, we have treated the target labels used in model training and calibration as reliable representations of the GT. In real-life applications, this might not be the case, and measures to deal with label noise might be needed. Second, the confidence scores are only point estimates of the empirical probabilities. A more thorough treatment would try to incorporate the uncertainty in these estimates as well. We leave the details of this treatment for future work.
Although this study has shown that the estimates produced by AC can compete with those provided by more complex methods and in many cases even beat them, the results and the theoretical analysis of AC should not be taken to imply that it is some magical bullet for all unsupervised monitoring problems. As stated by Garg et al. (2022), no estimator can prevail under concept shift and no estimator is the best in all cases. Our experiments provide an empirical demonstration of this. Without access to GT labels, there is no reliable way to know if concept shift is happening. Thus, confidence-based estimators should only be relied on in scenarios that can be expected to contain no concept shift.
@@ 6.3 Limitations and validity
In our experiments, we relied solely on synthetic datasets. Earlier studies on unsupervised accuracy estimation have focused on image classification with a minor interest also in textual data. In this work, we extend the scope of analysis to tabular data. Unfortunately, to the best of our knowledge, there are no real-life tabular datasets for covariate shift scenarios. Only recently, a benchmark for exploring distributional shifts in tabular data was published (Gardner et al., 2024), but all of the datasets in this benchmark express concept shift, rendering them unusable for our purposes.
Since the AC method hinges on the calibration assumption, it is worthwhile to ponder how reasonable this assumption is in reality. An often cited study (Ovadia et al., 2019) shows that the uncertainty estimates of most models deteriorate with covariate shift. However, the study in question deals only with multi-class image classification and equates calibration with Temperature Scaling (Guo et al., 2017). It is possible that the same conclusion might not hold for other scenarios or calibration techniques. In fact, promising techniques on how to remain calibrated under covariate shift are already starting to emerge (Białek et al., 2024). Our experiments show that if no interventions are made, the calibration error seems to increase slightly but steadily with increasing shift, also in a binary classification setting using Isotonic Regression (Zadrozny and Elkan, 2002). We also experimented with Beta Calibration (Kull et al., 2017) to derive the calibration mappings with similar results. However, although correlated, the estimation errors do not seem to increase as rapidly as the calibration errors do, implying that the performance of the estimators remains reasonable even if the calibration error did increase with shift.
Although the calibration errors for XGB and LGBM were almost unaltered by the increasing shift with both signaling low calibration errors even without applying a calibration mapping, this effect might be an artifact stemming from the simplicity of our synthetic datasets. In a more high-dimensional setting, gradient boosting can lead to high calibration error as well (Kivimäki et al., 2023). However, some models or ways to train models might be more robust against shift-induced miscalibration. We speculate that this might also provide a more credible explanation for why in (Guillory et al., 2021), the (regression-free) AC method performs better than DoC with Augmix-DeepAugment since training with Augmix is known to reduce miscalibration (Hendrycks* et al., 2020). Although perfect calibration is unattainable in real-life scenarios and calibration tends to deteriorate under shift, we encourage future research to explore further methods to help maintain sufficient calibration under covariate shift, enhancing the performance of confidence-based estimators.

### v2

@@ 4 Average Confidence
In this section, we present some theoretical results for the AC method, which was originally proposed as a baseline method for detecting misclassified and out-of-distribution instances in neural networks (?). When used in unsupervised accuracy estimation, the predictive accuracy is estimated to be the average of confidence scores, hence the name.
@@
Since $|F_{k}|=\binom{n}{k}$, calculating the required probabilities using the definition directly is infeasible even for relatively small values of $n$. Fortunately, there are several ways to avoid this combinatorial explosion. In this work, we leverage the algorithm presented by ? (?), which is based on the Fast Fourier Transform and implemented in the poibin [^1] https://github.com/tsakim/poibin python library. It allows us to derive the Poisson binomial distributions exactly and almost instantly.
@@ 4.3 Estimating the Confusion Matrix for Failure Prediction
Sometimes confidence scores are used as a tool for failure prediction (?, ?). A failure prediction mechanism based on confidence scores is a mapping $h:S\rightarrow\{0,1\}$. It sets a confidence threshold $t\in[0,1]$ and treats all predictions with a confidence score equal to or above the threshold as correct and below the threshold as incorrect. Predictions with sub-threshold confidence can be altogether discarded, or sent to further inspection by some downstream system or human-in-the-loop. This might be useful, especially in situations where different types of errors induce different costs. It is worthwhile to stress that this treatment creates second-order predictions (predictions about predictions) and is not to be confused with the true underlying discriminator $g$.
@@ 5 Experiments
We conduct two experiments [^2] Code is publicly available at https://github.com/JuhaniK/AC_trials with synthetic data. In the first experiment, we examine the quality of point predictions and confidence intervals of the estimates derived using the AC method when the confidence scores are known to be calibrated. In the second experiment, we compare AC, DoC-Feat, and ATC under covariate shift when used to estimate predictive accuracy in a monitoring setting.
@@ 5.1.1 Data Description
We chose to create our simulated set of confidence scores by drawing samples from a mixture of three Beta distributions. The first component was biased to generate high-confidence predictions, the second to generate average ones, and the third to create low-confidence predictions. By adjusting the mixture weights, we could effectively emulate distributions of confidence scores outputted by models operating on real-life data. Then, by changing the mixture weights, we could easily simulate shifts in model performance to see how the estimator under scrutiny managed to keep track of them. The original dataset was trying to emulate a typical distribution of confidence scores where most predictions are made with high confidence (?). For the shifted distribution, the fraction of high-confidence predictions was slightly decreased. The exact parameters and weights of the mixture components used in the experiment are given in Table 2.
@@
Table 2: Parameters and weights of the Beta mixture components
@@
Figure 2: Distribution of confidence scores in the simulated data.
@@ 5.1.2 Quality of Point Estimates under Shift
We tested the quality of point estimates for predictive accuracy derived using AC by simulating gradual data shifts. We used two different monitoring window sizes, namely 100 and 500 predictions. We gradually increased the fraction of instances sampled from the shifted distributions in increments of $5\%$, starting from $0\%$ and ending in $100\%$. For each degree of shift, we conducted 1,000 trials, where we measured the error of the estimate in each of the trials and plotted the mean of the errors in Fig. 3, along with one standard deviation (shaded area). We contrasted the results of AC by also similarly plotting estimates derived by using either ATC (?) or a binomial assumption (using the overall accuracy of the original distribution as the parameter). The binomial assumption comparison was included to give a sense of how fast the estimation error would increase if the predictive performance was assumed not to change.
@@
Figure 3: The quality of point estimates under gradual data drift.
@@ 5.1.3 Quality of Confidence Intervals
where $\sigma$ is an estimation of the population variance as the variance of the true labels in the reference data set (where GT is known) and $n$ is the monitoring window size (NannyML uses the term ”chunk size”). We replaced the ±$\pm 3$ standard errors with ±$\pm 1.96$ standard errors in our tests to get the lower and upper $95\%$ confidence limits for a fair comparison against the CIs derived using the Poisson binomial approach. The results of this experiment are presented in Fig. 4, clearly showing the superiority of the Poisson binomial approach, which produces (roughly) valid CIs. In contrast, the confidence bands produced by the SEM approach are extremely conservative.
@@
We argue that it is conceptually clearer to account for the variance due to sampling effects with control limits and form the confidence bands for the estimated accuracy solely on the basis of uncertainty in the estimates themselves using the Poisson binomial approach. In this way, all available information from each batch gets utilized and the width of the confidence band is allowed to vary from batch to batch. We are currently working to implement this improved approach to forming confidence bands into the NannyML library.
5.2 Experimenting under Covariate Shift
Next, we compared the confidence-based accuracy estimators AC, DoC-feat, and ACT with both uncalibrated and calibrated confidence scores. Earlier studies (?, ?) have had conflicting results on whether calibration has a beneficial effect in estimating predictive accuracy or not. In these studies, accuracy estimation has been performed on the dataset level (with sample size in the order of thousands or tens of thousands), whereas we are interested in estimates derived for sample sizes generally used in model monitoring (in the order of hundreds). Furthermore, they have emphasized neural networks used in image classification, with the nature of shifts in the shifted datasets not fully controlled. We expanded this scope by comparing the estimation methods using seven non-neural classifier models, namely Logistic Regression (LR), (Gaussian) Naïve Bayes (NB), K-Nearest Neighbors (KNN), Support Vector Machine (SVM) with RBF kernel, and Random Forest (RF), using implementations provided by the scikit-learn library  [^4] https://scikit-learn.org/stable/, as well as XGBoost (XGB)  [^5] https://xgboost.readthedocs.io/en/stable/, and Light-GBM (LGBM)  [^6] https://lightgbm.readthedocs.io/en/latest/index.html to see whether the earlier results would generalize to these models as well. All models were trained with the default parameter settings of their respective implementations.
@@ 5.2.1 Data and Model Training Descriptions
We used simulated data in order to have full control over the nature of the shift. Since all confidence-based accuracy estimators will eventually fail under concept shift (?), we restricted the shift as covariate shift. We experimented with two scenarios, where in both scenarios our data consisted of two features and a binary label (1 or 0).
@@
In both scenarios, we trained the models with 100,000 samples, of which $80\%$ were easy and $20\%$ were hard. Using the same ratio, we drew 25,000 additional samples to train the calibration mappings for each model and 25,000 additional samples for the setup required by the DOC-Feat and ATC methods. Since there was no shortage of data, we used the non-parametric Isotonic Regression (?) to derive the calibration mappings. It is generally considered to produce good quality calibration mappings but is also known to sometimes overfit with smaller datasets (?). In each trial, we created a test dataset of 25,000 samples, with an increasing portion of hard-to-predict samples, as described in Table 3.
@@
Table 3: The number of easy and hard samples for each degree of shift
@@ 5.2.2 Results with Linear Decision Boundary
Table 5: Mean absolute error of estimated accuracy using data with linear decision boundary. ”Acc” stands for Accuracy, ”ACE” for Adaptive Expected Calibration Error, ”AC” for Average Confidence, ”DOC” for DoC-Feat, and ”ATC” for Average Thresholded Confidence. Underscipts $u$ and $c$ refer to using uncalibrated or calibrated confidence scores respectively. All values are the averages over the 1,000 trials, except for calibration errors, which are calculated on the dataset level. The best estimator for each model and degree of shift is bolded.
@@ 5.2.3 Results with Non-linear Decision Boundary
Table 7: Mean absolute error of estimated accuracy using data with non-linear decision boundary. ”Acc” stands for Accuracy, ”ACE” for Adaptive Expected Calibration Error, ”AC” for Average Confidence, ”DOC” for DoC-Feat, and ”ATC” for Average Thresholded Confidence. Underscipts $u$ and $c$ refer to using uncalibrated or calibrated confidence scores respectively. All values are the averages over the 1,000 trials, except for calibration errors, which are calculated on the dataset level. The best estimator for each model and degree of shift is bolded.
@@ 6 Discussion
In this section, we analyze the significance of our findings. We underline some properties of confidence-based estimators in general and discuss the limitations and validity of our results.
@@ 6.1 Novel Contributions
Earlier studies (e.g., ?, ?, ?, ?, ?, ?) have explored the problem of unsupervised accuracy estimation only on the dataset level, that is, whether the suggested methods would be able to correctly estimate the accuracy of an ML model on some given shifted dataset, in the absence of GT labels. Furthermore, theoretical guarantees for the proposed methods are provided only by ? (?) and ? (?). In real-life model monitoring, instead of complete, shifted datasets, users typically have access only to relatively small samples of the potentially shifted data and are mostly interested in detecting relevant changes in the underlying data distribution on the fly to trigger appropriate interventions. In this study, we have tried to bridge this gap by exploring different confidence-based performance estimators in a monitoring setting with relatively small sample sizes.
@@ 6.2 Remarks on AC and Other Confidence-based Estimators
Ever since AC was introduced as a method for OoD-detection (?), it has become a de facto baseline for confidence-based estimators. However, up until now, there has never been a proper justification for its use as such or a thorough analysis of its properties. As confidence-based estimators (such as CBPE in NannyML) are already being implemented in monitoring libraries intended for industrial use, such theoretical analysis is long overdue. In this paper, we have provided such a justification by showing that under the calibration assumption, AC is an unbiased and consistent estimator of model accuracy. Another desirable property of AC is that one can leverage the properties of the Poisson binomial distribution to derive CIs for the estimates, which are shown to be valid under the same assumption. None of the other methods have been reported to have this property.
Contributions to the practical utility of AC by the theoretical and experimental findings presented in this paper are three-fold. First, they guarantee that if a model is known to be calibrated, a user can confidently rely on AC to produce good quality estimates of model accuracy on average. Second, the deviations in these estimates can be accounted for using CIs, which are also guaranteed to be valid under the calibration assumption. Third, the strong correlation between calibration and estimation errors for AC shows that using AC in unsupervised accuracy estimation reduces the problem of estimation into a problem of calibration, which is a widely studied and constantly developing research area.
Although using control limits and CIs to account for sampling effects and uncertainty in the estimates respectively, two other sources of uncertainty remain. First, in this work, we have treated the target labels used in model training and calibration as reliable representations of the GT. In real-life applications, this might not be the case, and measures to deal with label noise might be needed. Second, the confidence scores are only point estimates of the empirical probabilities. A more thorough treatment would account for the uncertainty in these estimates as well. We leave the details of this treatment for future work.
Although this study has shown that the estimates produced by AC can compete with those provided by more complex methods and in many cases even beat them, the results and the theoretical analysis of AC should not be taken to imply that it is some magical bullet for all unsupervised monitoring problems. As stated by ? (?), no estimator can prevail under concept shift and no estimator is the best in all cases. Our experiments provide an empirical demonstration of this. Without access to GT labels, there is no reliable way to know if concept shift is happening. Thus, confidence-based estimators should only be relied on in scenarios that can be expected to contain no concept shift.
@@ 6.3 Limitations and Validity
In our experiments, we relied solely on synthetic datasets. Earlier studies on unsupervised accuracy estimation have focused on image classification with a minor interest also in textual data. In this work, we extend the scope of analysis to tabular data. Unfortunately, to the best of our knowledge, there are no real-life tabular datasets for covariate shift scenarios. Only recently, a benchmark for exploring distributional shifts in tabular data was published (?), but all of the datasets in this benchmark express concept shift, rendering them unusable for our purposes.
Since the AC method hinges on the calibration assumption, it is worthwhile to ponder how reasonable this assumption is in reality. An often cited study (?) shows that the uncertainty estimates of most models deteriorate with covariate shift. However, the study in question deals only with multi-class image classification and equates calibration with Temperature Scaling (?). It is possible that the same conclusion might not hold for other scenarios or calibration techniques. In fact, promising techniques on how to remain calibrated under covariate shift are already starting to emerge (?). Our experiments show that if no interventions are made, the calibration error seems to increase slightly but steadily with increasing shift, also in a binary classification setting using Isotonic Regression (?). We also experimented with Beta Calibration (?) to derive the calibration mappings with similar results. However, although correlated, the estimation errors do not seem to increase as rapidly as the calibration errors do, implying that the performance of the estimators remains reasonable even if the calibration error did increase with shift.
Although the calibration errors for XGB and LGBM were almost unaltered by the increasing shift with both signaling low calibration errors even without applying a calibration mapping, this effect might be an artifact stemming from the simplicity of our synthetic datasets. In a more high-dimensional setting, gradient boosting can lead to high calibration error as well (?). However, some models or ways to train models might be more robust against shift-induced miscalibration. We speculate that this might also provide a more credible explanation for why in (?), the (regression-free) AC method performs better than DoC with Augmix-DeepAugment since training with Augmix is known to reduce miscalibration (?). Although perfect calibration is unattainable in real-life scenarios and calibration tends to deteriorate under shift, we encourage future research to explore further methods to help maintain sufficient calibration under covariate shift, enhancing the performance of confidence-based estimators.